solve: 1/3^2-(-1/2^3)(1/3-(-1/2)
Answer by stanbon(75887)

1/3^2-(-1/2^3)(1/3-(-1/2)
= (1/9)-(1/8)((1/3)+1/2))
= (1/9)-(1/8)(5/6)
= (1/9)-(5/48)
= (3/[9*48]
= (1/144)
===============
Cheers,
Stan H.